# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

### v1

@@ IV General Lower Bound
    $\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
    $=\min_{a\notin\{1,2,\cdots,M\}}\bigg{(}\sum_{i=1}^{M}w_{i}+w_{a}\bigg{)}I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\frac{w_{2}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a}).$ (32)
@@
$T^{*}(\mu)^{-1}$ $=\sup_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (33)
@@
$w^{*}(\mu)=\text{argmax}_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (34)
@@ V-A Sampling Rule: Tracking the Optimal Proportions
We adopt the same sampling rules as described in (Garivier2016OptimalBA, , Section 3). The core idea is to match the optimal sampling proportions $w^{*}(\mu)$ by tracking the plug-in estimate $w^{*}(\hat{\mu}(t))$. However, in bandit settings, relying directly on plug-in estimates can be risky: poor initial estimates may cause some arms to be prematurely discarded, thus preventing the collection of new observations that would correct earlier mistakes. In fact, naive plug-in tracking may fail to identify the best arm reliably.
@@
$A_{t+1}\in\begin{cases}\arg\min_{a\in U_{t}}N_{a}(t)&\text{if }U_{t}\neq\emptyset\quad\text{(\emph{forced exploration})},\\ \arg\max_{1\leq a\leq K}\left[tw^{*}_{a}(\hat{\mu}(t))-N_{a}(t)\right]&\text{otherwise (\emph{direct tracking})}.\end{cases}$ (38)
Both sampling rules are guaranteed to ensure that the empirical sampling proportions converge to the optimal allocation, as established in (Garivier2016OptimalBA, , Section 3.1).
@@ V-B Chernoff’s Stopping Rule
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)=\log\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})},$ (39)
@@
$p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})$ $=\prod_{k=1}^{N_{a}(t)}\exp\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{a})X_{k}-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg{)}$ (41)
$=\exp\bigg{[}N_{a}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{a})\hat{\mu}_{a}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg{)}\bigg{]}.$ (42)
@@
$p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$ $=\exp\bigg{[}N_{b_{i}}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{b_{i}})\hat{\mu}_{b_{i}}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{b_{i}}))\bigg{)}\bigg{]},\qquad\forall i\in[M].$ (44)
Hence, we have
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]},$ (45)
@@
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg{)},\qquad\forall x\in\{a,b_{1},b_{2},\cdots,b_{M}\},$ (46)
is concave in $\mu^{\prime}_{x}$.
@@
$=-N_{x}(t)(\hat{\mu}_{x}(t)-\mu^{\prime}_{x})\bigg{(}\frac{\ddot{A}(\mu^{\prime}_{x})}{(\dot{A}(\mu^{\prime}_{x}))^{2}}\bigg{)},$ (49)
@@
$\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}\geq l_{a}(\hat{\mu}_{a}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (50)
@@
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
@@
$\log\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]},$ (52)
@@
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg{)},\qquad\forall x\in\{a,b_{1},b_{2},\cdots,b_{M}\},$ (53)
is concave in $\mu^{\prime}_{x}$.
@@
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}.$ (54)
@@
$L(\mu^{\prime}_{a},\mu^{\prime}_{b_{1}},\mu^{\prime}_{b_{2}},\cdots,\mu^{\prime}_{b_{M}},\beta_{1},\beta_{2},\cdots,\beta_{M})=l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})+\sum_{i=1}^{M}\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}})$ (55)
@@
00 0 $=\frac{\partial L}{\partial\mu^{\prime}_{a}}=N_{a}(t)\big{(}\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{a})}{(\dot{A}(\mu^{\prime}_{a}))^{2}}\bigg{]}+\sum_{i=1}^{M}\beta_{i},$ (56)
00 0 $=\frac{\partial L}{\partial\mu^{\prime}_{b_{i}}}=N_{b_{i}}(t)\big{(}\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{b_{i}})}{(\dot{A}(\mu^{\prime}_{b_{i}}))^{2}}\bigg{]}-\beta_{i},\qquad\forall i\in[M],$ (57)
00 0 $=\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}}),\qquad\forall i\in[M]$ (58)
@@
Since $\beta_{i}\geq 0$ for all $i\in[M]$ from (56) and the fact that $\ddot{A}(\mu^{\prime}_{a})>0$ (exponential bandits), we have $\hat{\mu}_{a}(t)\geq\mu^{\prime}_{a}$. Now, from (58) if there exists $i\in[M]$ such that $\beta_{i}=0$, then from (57) we also have $\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t)$. Hence, it follows from (59) we have
$\hat{\mu}_{a}(t)\geq\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t).$ (60)
Under the condition that $\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, it holds from (60) that
$\hat{\mu}_{a}(t)=\mu^{\prime}_{a}=\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t).$ (61)
From (61) and (58), under the condition that $\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, we must have
@@
$N_{a}(t)\big{(}\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{a})}{(\dot{A}(\mu^{\prime}_{a}))^{2}}\bigg{]}+\sum_{i=1}^{M}N_{b_{i}}(t)\big{(}\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{b_{i}})}{(\dot{A}(\mu^{\prime}_{b_{i}}))^{2}}\bigg{]}=0.$ (63)
From (62) and (63), under the condition that $\hat{\mu}_{a}(t)\geq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, we obtain
$\mu^{\prime}_{a}=\mu^{\prime}_{b_{1}}=\mu^{\prime}_{b_{2}}=\cdots=\mu^{\prime}_{b_{M}}=\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t):=\frac{N_{a}(t)\hat{\mu}_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{N_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)}.$ (64)
@@
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}=l_{a}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$ (65)

### v2

@@ IV General Lower Bound
    $\min_{a\notin[M]}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}$
    $\qquad\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
    $=\min_{a\notin[M]}\bigg(\sum_{i=1}^{M}w_{i}+w_{a}\bigg)$
    $\qquad\times I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\cdots,\mu_{M},\mu_{a}).$ (38)
In cases where a subset of $\{w_{1},w_{2},\ldots,w_{M}\}$ takes zero values, it is straightforward to verify that the optimal value of the problem remains unchanged from that in (38).
@@
$T^{*}(\mu)^{-1}$ $=\sup_{w\in\Sigma_{K}}\min_{a\notin[M]}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}$
$\qquad\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (39)
@@
$w^{*}(\mu)$ $=\text{argmax}_{w\in\Sigma_{K}}\min_{a\notin[M]}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}$
$\qquad\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (40)
@@ V-A Sampling Rule: Tracking the Optimal Proportions
We adopt the same sampling rules as described in [3, Section 3]. The core idea is to match the optimal sampling proportions $w^{*}(\mu)$ by tracking the plug-in estimate $w^{*}(\hat{\mu}(t))$. However, in bandit settings, relying directly on plug-in estimates can be risky: poor initial estimates may cause some arms to be prematurely discarded, thus preventing the collection of new observations that would correct earlier mistakes. In fact, naive plug-in tracking may fail to identify the best arm reliably.
@@
$A_{t+1}\in\begin{cases}\arg\min_{a\in U_{t}}N_{a}(t)&\text{if }U_{t}\neq\emptyset,\\ \arg\max_{1\leq a\leq K}\left[tw^{*}_{a}(\hat{\mu}(t))-N_{a}(t)\right]&\text{otherwise}.\end{cases}$ (44)
Both sampling rules are guaranteed to ensure that the empirical sampling proportions converge to the optimal allocation, as established in [3, Section 3.1].
@@ V-B Chernoff’s Stopping Rule
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)=\log\bigg($
$\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\bigg),$ (45)
@@
$p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})$
$\quad=\prod_{k=1}^{N_{a}(t)}\exp\bigg(\dot{A}^{-1}(\mu^{\prime}_{a})X_{k}-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg)$ (47)
$\quad=\exp\bigg[N_{a}(t)\bigg(\dot{A}^{-1}(\mu^{\prime}_{a})\hat{\mu}_{a}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg)\bigg].$ (48)
@@
$p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\quad=\exp\bigg[N_{b_{i}}(t)\bigg(\dot{A}^{-1}(\mu^{\prime}_{b_{i}})\hat{\mu}_{b_{i}}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{b_{i}}))\bigg)\bigg]$ (50)
for all $i\in[M]$. Hence, we have
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg],$ (51)
@@
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg(\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg),$ (52)
for all $x\in\{a,b_{1},b_{2},\cdots,b_{M}\}$, is concave in $\mu^{\prime}_{x}$.
@@
$=N_{x}(t)(\hat{\mu}_{x}(t)-\mu^{\prime}_{x})\bigg(\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{x})}\bigg),$ (55)
@@
$\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg]$
$\quad\leq l_{a}(\hat{\mu}_{a}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (56)
@@
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
@@
$\log\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg],$ (58)
@@
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg(\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg),$ (59)
for all $x\in\{a,b_{1},b_{2},\cdots,b_{M}\}$, is concave in $\mu^{\prime}_{x}$.
@@
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg].$ (60)
@@
$L(\mu^{\prime}_{a},\mu^{\prime}_{b_{1}},\mu^{\prime}_{b_{2}},\cdots,\mu^{\prime}_{b_{M}},\beta_{1},\beta_{2},\cdots,\beta_{M})$
$\quad=l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})+\sum_{i=1}^{M}\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}})$ (61)
@@
$0$ $=\frac{\partial L}{\partial\mu^{\prime}_{a}}=N_{a}(t)\big(\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big)\bigg[\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{a}))}\bigg]+\sum_{i=1}^{M}\beta_{i},$ (62)
$0$ $=\frac{\partial L}{\partial\mu^{\prime}_{b_{i}}}=N_{b_{i}}(t)\big(\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big)\bigg[\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{b_{i}}))}\bigg]-\beta_{i},$ (63)
$0$ $=\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}}),\qquad\forall i\in[M]$ (64)
@@
Now, let $\mathcal{U}=\{i\in[M]:\beta_{i}\neq 0\}$ and $\mathcal{V}=\{i\in[M]:\beta_{i}=0\}$. We consider two cases:
  • Case 1: $\mathcal{V}=\emptyset$. Then, we have $\mathcal{U}=[M]$. From (64), we must have
@@
    $N_{a}(t)\big(\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big)\bigg[\frac{1}{\ddot{A}(A^{-1}(\mu^{\prime}_{a}))}\bigg]$
    $\quad+\sum_{i=1}^{M}N_{b_{i}}(t)\big(\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big)\bigg[\frac{1}{\ddot{A}(A^{-1}(\mu^{\prime}_{b_{i}}))}\bigg]=0.$ (67)
    From (66) and (67), we obtain
    $\mu^{\prime}_{a}=\mu^{\prime}_{b_{1}}=\mu^{\prime}_{b_{2}}=\cdots=\mu^{\prime}_{b_{M}}=\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)$
    $\quad:=\frac{N_{a}(t)\hat{\mu}_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{N_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)}.$ (68)
@@
    $\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg]$
    $\quad=l_{a}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$ (69)